# Design Patterns for the Common Good: Building Better Technologies Using the Wisdom of Virtue Ethics

### v1

@@ 7.5. Limitations and Opportunities for Future Research
The design patterns we created for social media were a proof of concept for a virtue-based design process. Because they were inspired by Catholic Social Teaching, they reflect one particular set of virtues in their design. Our participants generally liked the designs but wanted there to be additional options, especially given that our designs prioritized intentional conversation at the expense of information-gathering. This leaves room for future work to study what connecting widely and information-gathering would look like when done virtuously, especially minimizing exploitative mechanisms and impersonal or disrespectful conversations. One way to foster creativity in this process is to repeat it with virtues from other traditions. Are there insights we can gain from considering Confucian, Buddhist, Hindu, or Islamic virtues that are lost in the western perspective that could help us to reconcile the dichotomy between intentional conversation and exposure to new ideas? Considering other virtue traditions would also address the call for more ethical pluralism in HCI (Ahmed, 2022; Rifat et al., 2023).
@@
Finally, one limitation of our approach is that other aspects of the technology development process, including design processes and company organization, are left out when focusing only on design patterns. We did not consider the Catholic Social Teaching principles the dignity of work and rights of workers and rights and responsibilities in our approach because they pertain more to how the company is run than to the design of the technology itself. Additionally, the option for the poor and vulnerable included “prioritize the needs of marginalized groups in design” which could entail incorporating the community-based design practices advocated by the Design Justice (Costanza-Chock, 2020) movement into the design process. We recommend future work considering how virtue can be incorporated into other aspects of how companies are run.

### v2

@@ 7.5. Limitations and Opportunities for Future Research
The design patterns we documented for social media were a proof of concept for a virtue-based design process. Because they were inspired by Catholic Social Teaching, they reflect one particular set of virtues in their design. Our participants generally liked the designs but wanted there to be additional options, especially given that our designs prioritized intentional conversation at the expense of information-gathering. This leaves room for future work to study what connecting widely and information-gathering would look like when done virtuously, especially minimizing exploitative mechanisms and impersonal or disrespectful conversations. One way to foster creativity in this process is to repeat it with virtues from other traditions. Are there insights we can gain from considering Confucian, Buddhist, Hindu, or Islamic virtues that are lost in the western perspective that could help us to reconcile the dichotomy between intentional conversation and exposure to new ideas? Considering other virtue traditions would also address the call for more ethical pluralism in HCI (Ahmed, 2022; Rifat et al., 2023).
@@
Finally, one limitation of our approach is that other aspects of the technology development process, including design processes and company organization, are left out when focusing only on design patterns. We did not consider the Catholic Social Teaching principles dignity of work and rights of workers and rights and responsibilities in our approach because they pertain more to how the company is run than to the design of the technology itself. Additionally, the option for the poor and vulnerable included “prioritize the needs of marginalized groups in design” which could entail incorporating the community-based design practices advocated by the Design Justice (Costanza-Chock, 2020) movement into the design process. We recommend future work considering how virtue can be incorporated into other aspects of how companies are run.
@@ 8. Conclusion
Acknowledgements.
This work was supported by the Responsible Computing Challenge, a partnership of Omidyar Network, Mozilla, Schmidt Futures, Craig Newmark Philanthropies and Mellon Foundation. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of our sponsors.
Appendix A Appendix
Illustrations of the technical inquiry process for the other six design patterns whose illustrations were not included in the main body of this paper are found on the following pages.
Figure 10. Technical Inquiry process for the Friends Over Followers pattern.
[…]
On the top there is a green bubble. It includes the following text: App: BeReal. Design: Friend connection structure and message on the add friends page encouraging people to only add people they know in real life. CST Upholding: Friend connection structures encourage connecting in smaller groups, upholding subsidiarity. When a user has fewer connections, interactions are more private and personal, promoting life and dignity of the human person.On the bottom there is a red bubble. It includes the following text: App: X, TikTok. Design: Follower connection structure. CST Violation: Follower connection structures allow for unlimited and less personal interactions, violating subsidiarity The two bubbles have arrows to a blue box that says: Design Pattern: Friends Over Followers. Intent: Encourage that free-form, frequent, and more vulnerable online sharing occur only with closer and fewer connections. Solution: The more free-forming the sharing is, the more that should be done in small groups with “friend” connections where both parties have to add each other in order to interact. Free-forming sharing can include long discussions without moderation, the sharing of personal images that could be filtered or doctored, etc. This pattern can optionally include a prompt to add only people one knows in real life, saying something along the lines of, “Are you sure you know them well?” (This specific example is taken from BeReal.) Consequences: If this pattern was employed all the time, then some good uses of “following,” like thoughtful Substack blogs, would be eliminated. Virtues Embodied: Life and Dignity of the Human Person, Subsidiarity
Figure 11. Technical Inquiry process for the Moderated Mingling pattern.
[…]
On the top there are two green bubbles and on the bottom there is one green bubble and one red bubble. The top left green bubble has the following text: App: Discord, X. Design: Groups for particular topics or interests. CST Upholding: This upholds subsidiarity by keeping groups topic-specific. The top right bubble has the following text: App: Discord, X. Design: Ability to join large group conversations or communities with a wide variety of participants. CST Upholding: This promotes solidarity by facilitating connections with the greater human family. The bottom left green bubble has the following text: App: Discord, X. Design: Groups are moderated by community members. CST Upholding: This upholds subsidiarity through keeping the moderation self-directed. The bottom right red bubble has the following text: App: X, TikTok. Design: Feed-style sharing with anyone on the app. CST Violation: The ability to see content from all users in an impersonal feed format can lead to less intentional and in some cases hostile interactions, violating life and dignity of the human person and option for the poor and vulnerable. The bubbles have arrows to a blue box that says: Design Pattern: Moderated Mingling. Intent: For larger-group communities, have focused intentions for the conversation and keep conduct respectful. Solution: Encourage self-moderated communities where communities set their own terms of moderation, especially focused on keeping the conversation on topic and the conduct respectful. Consequences: Users may feel their freedom is violated by having to follow moderation guidelines. Virtues Embodied: Life and Dignity of the Human Person, Option for the Poor and Vulnerable, Solidarity, Subsidiarity.
Figure 12. Technical Inquiry process for the Clear Algorithmic Comprehension pattern.
[…]
There is a red bubble with the following text: App: TikTok, X. Design: Attention-based and non transparent content recommendation algorithms. CST Violation: This violates option for the poor and vulnerable by employing attention-based profit models and subsidiarity by taking control of what one sees away from the user, thus having decisions be made at a higher level than appropriate. There is an arrow from the red bubble to a blue box with the following text: Design Pattern: Clear Algorithmic Comprehension. Intent: Allow users to know why types of content are being recommended to them and give them greater control over recommended content. Solution: Make content recommendation algorithms as transparent as possible to reveal why they recommend the content that they do. Furthermore, make algorithms customizable so users can add and remove the types of content that they want to be shown. Additionally include the capacity to turn off recommended content. Consequences: This is technically difficult to do. Virtues Embodied: Option for the Poor and Vulnerable, Subsidiarity.
Figure 13. Technical Inquiry process for the Moderated Entry pattern.
[…]
There are two green bubbles. The green bubble on the left has the following text: App: Discord. Design: Groups may set entry criteria for participants. CST Upholding: Having entry critera for participants gives greater control over group members, upholding subsidiarity. When group members are vetted to ensure they are real people who will contribute respectfully, this upholds life and dignity of the human person. The green bubble on the right has the following text: App: X (Premium Version). Design: Verified interactions through government ID. CST Upholding: This promotes respectful conversation when people show up with their real identity, upholding life and dignity of the human person. There is an arrow pointing from the green bubbles to a blue box with the following text: Design Pattern: Moderated Entry. Intent: For large-group forums or discussions, restrict entry to make sure conversations stay on topic, are respectful and that real people (as opposed to spam bots) are engaging in the conversation. Solution: Have an identity verification process or some sort of other vouching process (e.g. via phone number, email address, personal conversation, etc.) to ensure group members are real people who will make meaningful contributions. Consequences: In some cases, anonymity is desirable. For example, if those in marginalized communities want to communicate without retaliation. In these cases, employing moderation, smaller-group connecting, or singularly-focused applications are especially important. To avoid cliquishness, this design pattern should be balanced out with other patterns to encourage enough openness for productive dialogue. Moderation criteria should not be unreasonably exclusive. Virtues Embodied: Life and Dignity of the Human Person, Subsidiarity
Figure 14. Technical Inquiry process for the No Lurking pattern.
[…]
There is a green bubble on top of a red bubble. The green bubble has the following text: App: BeReal. Design: Posts of others are hidden until the user makes a post of their own. CST Upholding: Regular particpation and communication can help to strengthen relationships, upholding the call to family, community, and participation. The red bubble has the following text: App: X, TikTok. Design: Feed-style sharing. CST Violation: Feed-style sharing can encourage passive consumption over active participation, violating the call to family, community, and participation. There are arrows coming from the bubbles to a blue box with the following text: Design Pattern: No Lurking. Intent: Encouraging participation in conversations rather than just lurking without participating. Solution: Hide the posts of others until the user themselves makes an initial post. Consequences: Depending on the purpose of the application, lurking may be acceptable. The more the purpose of the application entails intentional participation from every member the more this pattern should be considered. Virtues Embodied: Call to Family, Community, and Participation
Figure 15. Technical Inquiry process for the Notification Intentionality pattern.
[…]
There is a red bubble with the following text: App: BeReal. Design: Incessant notifications encouraging you to open the app to see what your friends posted. CST Violation: This violates life and dignity of the human person and option for the poor and vulnerable, and care of God’s creation by demanding more attention than necessary to use the app well. There is an arrow pointing from the red bubble to a blue box with the following text: Design Pattern: Notification Intentionality. Intent: Notifications should encourage the user towards using the app in ways that promote connecting with others without being excessive. Solution: Applications should only send notifications that are essential for users to use the platform well. Any excessive notifications encouraging the user to open the app unnecessarily should be avoided. Consequences: Limited notifications could lead users to miss an ongoing conversation among other friends. Virtues Embodied: Life and Dignity of the Human Person, Option for the Poor and Vulnerable, Care of God’s Creation